### v2

@@ 2.5 Test-Time Adaptation
where $B=|\mathcal{D}_{\mathrm{tar}}|$ and $0<\rho<0.5$. Pseudo-labels are then defined by $\hat{y}_{t}=1$ if $t\in A$, and $\hat{y}_{t}=0$ if $t\in N$. Adaptation is performed by optimizing an image-level binary cross-entropy loss on the selected samples, i.e., $\mathcal{L}_{\mathrm{image}}=\mathrm{BCE}(\hat{y}_{i},\tilde{y}_{i})$ for $i\in A\cup N$.
@@ 3 Experiments
Benchmarks. To demonstrate the effectiveness of our method, we conduct experiments on 13 widely used benchmarks spanning industrial inspection and medical imaging. Specifically, we evaluate on six industrial datasets: BTAD [Mishra et al.(2021)Mishra, Verk, Fornasier, Piciarelli, and Foresti], VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer], MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger], DAGM [Wieler and Hahn(2007)], SDD [Tabernik et al.(2020)Tabernik, Sela, Skvarc, and Skocaj], and DTD-Synthetic [Aota et al.(2023b)Aota, Tong, and Okatani] and seven medical datasets, subdivided into detection-only (HeadCT [Kitamura(2018)], BrainMRI [Kanade and Gumaste(2015)], Br35H [Hamada(2020)]) and localization (ISIC [Codella et al.(2018)Codella, Gutman, Celebi, Helba, Marchetti, Dusza, Kalloo, Liopyris, Mishra, Kittler, et al.], CVC ColonDB [Tajbakhsh et al.(2015)Tajbakhsh, Gurudu, and Liang], CVC ClinicDB [Bernal et al.(2015)Bernal, Sánchez, Fernández-Esparrach, Gil, Rodríguez, and Vilariño], Kvasir [Jha et al.(2020)Jha, Smedsrud, Riegler, Halvorsen, De Lange, Johansen, and Johansen]). The medical datasets for anomaly localization contain only anomalous samples and lack normal data; since our TTA strategy requires normal samples, results are reported without TTA (see Appendix H). To ensure strict zero-shot evaluation, auxiliary datasets exclude all target data. More specifically, for a fair comparison, we followed the related literature [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi], and used MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger] as the auxiliary dataset when reporting the performance on all other datasets, while when reporting the performance on MVTec AD, we trained the model on VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer] as the auxiliary dataset.
Evaluation Metrics. Following standard practices in zero-shot anomaly detection literature such as  [Chen et al.(2023)Chen, Han, and Zhang, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], we evaluated our method using the AUROC. Additionally, we report APs for anomaly detection and use AUPRO [Bergmann et al.(2020)Bergmann, Fauser, Sattlegger, and Steger] to assess pixel-level anomaly localization performance. Results are provided at both dataset-level and domain-level averages.
Table 1: Model performance on public benmarks. The top section reports image-level results (AUROC, AP), and the bottom section reports pixel-level results (AUROC, AUPRO). Best entries are marked with bold and second-best entries with underline.
@@
Table 2: Comparison of PILOT vs. state-of-the-art on industrial benchmarks. $\mathbf{Img}$ and $\mathbf{Pix}$ denote image-level and pixel-level metrics, respectively. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. Arrows indicate change in performance: $\uparrow$(increase >0.5>0.5), $\downarrow$(decrease <0.5<-0.5), and $\circ$(no significant change, $\pm 0.5$).
@@
3.1 Comparison with State-of-the-Art
As reported in Table 1, PILOT establishes state-of-the-art performance across all benchmarks for both image-level anomaly detection and pixel-level localization in the zero-shot setting. This can be benefited from PILOT comprehensive approach in addressing critical challenges prevalent in ZSAD, that explicitly tackles domain shift in test phase. Prior approaches [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] rely solely on predefined prompts, which may not generalize well to unseen domains, whereas methods such as [Chen et al.(2023)Chen, Han, and Zhang, Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi, Zhou et al.(2022a)Zhou, Yang, Loy, and Liu, Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] employ lightweight visual adapters or a single learnable prompt to capture anomaly cues, but often overfit to the auxiliary training set, limiting robustness under distribution shift. PILOT distinguishes itself by dynamically identifying the most relevant prompts for each input and refining only those learnable prompt parameters during inference using high-confidence pseudo-labels from the unlabeled target data. Note that while existing baselines can utilize the proposed image-level TTA strategies, these often degrade pixel-level localization performance and do not yield consistent improvements across both detection and localization tasks. On the contrary, PILOT uniquely achieves gains in both image-level and pixel-level metrics under TTA, a phenomenon analyzed in detail in Section 3.2. We also compare against state-of-the-art full-shot methods in Appendix F.1, which show that PILOT achieves image-level accuracy on par.
@@ Appendix A Related Work
Zero-shot Anomaly Detection. ZSAD has gained increasing attention as a solution for identifying anomalies in unseen categories without requiring task-specific training data [Aota et al.(2023a)Aota, Tong, and Okatani, Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer]. Leveraging the generalization capabilities of Vision–Language Models (VLMs) such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], many approaches extend these models by linking image features to textual descriptions of normal and anomalous states. For instance, CLIP-AD [Chen et al.(2024)Chen, Zhang, Tian, He, Zhang, Wang, Wang, and Liu] and ZOC [Esmaeilpour et al.(2022)Esmaeilpour, Liu, Robertson, and Shu] are early studies that leverage CLIP for ZSAD, but primarily focus on anomaly classification rather than localization. WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer] uses multiple handcrafted prompts with iterative patch-based forward passes, and April-GAN [Chen et al.(2023)Chen, Han, and Zhang] refines local semantics via fine-tuning on an auxiliary dataset; both methods address image-level anomaly detection and pixel-level localization. More recent prompt learning methods, such as AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], AnomalyGPT [Gu et al.(2024)Gu, Zhu, Zhu, Chen, Tang, and Wang], and AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi], optimize object-agnostic text prompts in a single pass using auxiliary dataset. However, reliance on auxiliary datasets may introduce bias and compromise the intended zero-shot setting. This trade-off has been examined by Marzullo et al. [Marzullo and Ranzini(2024)], who observed that while avoiding auxiliary data reduces fine-tuning overhead, static prompts often fail to adapt to domain shifts; conversely, fine-tuning on auxiliary sets can improve accuracy but risks dataset bias and reduced generalization. Our work aims at addressing these challenges by introducing a dual-branch prompt learning mechanism that combines learnable and fixed prompts, enabling more robust and adaptive anomaly detection under domain shift.
@@
CLIP Adaptation. Recent advances in CLIP fine-tuning primarily rely on prompt learning, in which learnable tokens are introduced into the text encoder to align visual features with related semantics [Zhou et al.(2022b)Zhou, Yang, Loy, and Liu]. Approaches such as CoCoOp [Zhou et al.(2022a)Zhou, Yang, Loy, and Liu] and visual prompt tuning [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim] condition the prompts on image content to improve robustness under domain shift. MaPLe [Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan] extends prompt learning by inserting learnable tokens into both the visual and text encoders of CLIP, jointly optimizing them to enhance cross-modal alignment. These techniques achieve parameter-efficient adaptation by updating only the prompt representations and, in some cases, only CLIP’s final visual projection layer [Fahes et al.(2024)Fahes, Vu, Bursuc, Pérez, and de Charette]. Following this rationale, our method fine-tunes the visual projection layer to minimize additional parameters.
@@
Test Time Adaptation. A related direction for adapting CLIP-based models, particularly under domain shift, is TTA. In TTA, the model updates its parameters using unlabeled test data, often guided by pseudo-labels or self-supervised objectives. For instance, self-supervised losses based on the model’s own outputs can fine-tune representations without ground-truth labels [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt], and momentum-based pseudo-labeling helps enforce consistency across different augmented views [Goyal et al.(2022)Goyal, Sun, Raghunathan, and Kolter]. Other TTA strategies include entropy minimization, which guides the model to assign higher confidence to its predictions on unlabeled data, and the online updating of batch normalization statistics, which helps the model better align with the target distribution during test time. These approaches have been shown to improve robustness and generalization in image classification [Nado et al.(2020)Nado, Padhy, Sculley, D’Amour, Lakshminarayanan, and Snoek] and segmentation [Colomer et al.(2023)Colomer, Dovesi, Panagiotakopoulos, Carvalho, Härenstam-Nielsen, Azizpour, Kjellström, Cremers, and Poggi]. However, recent studies have found that these techniques are often less effective when the domain shift is large or when the model relies on strong vision-language priors, as in the case of CLIP. To our knowledge, AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] is the only TTA method for zero-shot anomaly localization, which employs an adapter module for pixel-level refinement and thus introduces additional parameters while focusing solely on localization. In contrast, our method adaptively updates the relevant learnable prompts, incurring no additional parameters and delivering consistent improvements in both image-level anomaly detection and pixel-level anomaly localization.
@@ B.1 Benchmark Datasets
Appendix B provides a comprehensive review of the thirteen benchmark datasets utilized in our study, as summarized in Table 4. To rigorously evaluate both image-level detection and pixel-level localization capabilities of PILOT, we select six industrial inspection datasets. DAGM [Wieler and Hahn(2007)] and DTD-Synthetic [Aota et al.(2023b)Aota, Tong, and Okatani] comprise textured surfaces exhibiting subtle variations in material appearance, presenting unique challenges for texture-based anomaly detection due to the fine-grained and ambiguous nature of defects. In contrast, MVTec AD [Bergmann et al.(2019)Bergmann, Fauser, Sattlegger, and Steger], VisA [Zou et al.(2022)Zou, Jeong, Pemula, Zhang, and Dabeer], BTAD [Mishra et al.(2021)Mishra, Verk, Fornasier, Piciarelli, and Foresti], and SDD [Tabernik et al.(2020)Tabernik, Sela, Skvarc, and Skocaj] focus on object-centric industrial scenarios, with defects such as cracks, scratches, dents, and missing components distributed across everyday manufactured and synthetic objects. These datasets are characterized by diverse object categories and high variability in anomaly manifestation, supporting comprehensive benchmarking of localization and generalization performance.
@@
To assess clinical applicability, we include seven medical imaging benchmarks. Three brain scan datasets, HeadCT [Kitamura(2018)], BrainMRI [Kanade and Gumaste(2015)], and Br35H [Hamada(2020)], comprise both healthy and pathological subjects, facilitating binary anomaly classification tasks on volumetric neuroimaging data. These datasets encompass a spectrum of anatomical and pathological variations, thereby enabling evaluation of robustness to inter-patient heterogeneity and rare conditions. The remaining four datasets provide pixel-wise lesion annotations and consist exclusively of abnormal samples. ISIC [Codella et al.(2018)Codella, Gutman, Celebi, Helba, Marchetti, Dusza, Kalloo, Liopyris, Mishra, Kittler, et al.] features dermoscopic images annotated for skin lesions, while CVC-ClinicDB [Bernal et al.(2015)Bernal, Sánchez, Fernández-Esparrach, Gil, Rodríguez, and Vilariño], CVC-ColonDB [Tajbakhsh et al.(2015)Tajbakhsh, Gurudu, and Liang], and Kvasir [Jha et al.(2020)Jha, Smedsrud, Riegler, Halvorsen, De Lange, Johansen, and Johansen] contain endoscopic images of the gastrointestinal tract, highlighting a range of polyp and ulcer appearances. Since these medical localization datasets lack normal examples, test-time adaptation is not performed for them; their localization results without TTA are reported in Appendix G.
@@ B.2 Domain Discrepancy Analysis
To validate that MVTec AD and VisA indeed represent statistically distinct domains suitable for strict zero-shot evaluation, we quantitatively assessed the distributional discrepancy between their CLIP global feature embeddings. Specifically, we employed Maximum Mean Discrepancy (MMD2[Gretton et al.(2012)Gretton, Borgwardt, Rasch, Schölkopf, and Smola], Fréchet Inception Distance (FID) [Heusel et al.(2017)Heusel, Ramsauer, Unterthiner, Nessler, and Hochreiter], and the Kolmogorov–Smirnov (KS) [Massey Jr(1951)] test. Our analysis yielded $\mathrm{MMD}^{2}=0.0328$ ($p$-value =0.0099=0.0099), indicating a significant difference in embedding distributions. The FID score between the two datasets is $51.9$, further corroborating a substantial distributional gap. Additionally, the KS test computed across all feature dimensions revealed that the minimum, mean, and maximum $p$-values are 0, $5.16\times 10^{-5}$, and $0.0329$, respectively, suggesting statistically significant differences in nearly every dimension of the embedding space. These results collectively confirm that MVTec AD and VisA are non-overlapping in the learned representation space, thus justifying their use as auxiliary and target datasets in our zero-shot anomaly detection experiments.

### v3

@@ 2.5 Test-Time Adaptation
where $B=|\mathcal{D}_{\mathrm{tar}}|$ and $0<\rho<0.5$. Pseudo‐labels are then defined by $\hat{y}_{t}=1$ if $t\in A$, and $\hat{y}_{t}=0$ if $t\in N$. Adaptation is performed by optimizing an image-level binary cross-entropy loss on the selected samples, i.e., $\mathcal{L}_{\mathrm{image}}=\mathrm{BCE}(\hat{y}_{i},\tilde{y}_{i})$ for $i\in A\cup N$.
@@ 3 Experiments
Benchmarks. To demonstrate the effectiveness of our method, we conduct experiments on 13 widely used benchmarks spanning industrial inspection and medical imaging. Specifically, we evaluate on six industrial datasets: BTAD [dataset_btad], VisA [dataset_visa], MVTec AD [dataset_mvtec], DAGM [dataset_DAGM], SDD [dataset_ksdd], and DTD‑Synthetic [dataset_DTD] and seven medical datasets, subdivided into detection-only (HeadCT [dataset_headCT], BrainMRI [dataset_brain], Br35H [dataset_br35h]) and localization (ISIC [dataset_isic], CVC ColonDB [dataset_ColonDB], CVC ClinicDB [dataset_ClinicDB], Kvasir [dataset_Kvasir]). The medical datasets for anomaly localization contain only anomalous samples and lack normal data; since our TTA strategy requires normal samples, results are reported without TTA (see Appendix H). To ensure strict zero-shot evaluation, auxiliary datasets exclude all target data. More specifically, for a fair comparison, we followed the related literature [AnomalyCLIP, AdaCLIP], and used MVTec AD [dataset_mvtec] as the auxiliary dataset when reporting the performance on all other datasets, while when reporting the performance on MVTec AD, we trained the model on VisA [dataset_visa] as the auxiliary dataset.
Baseline. We compare PILOT with the most recent SOTA methods, including SAA[saa], CoOp [CoCoOp], WinCLIP [winclip], APRIL-GAN [April_GAN], AnomalyCLIP [AnomalyCLIP], and AdaCLIP [AdaCLIP].
Evaluation Metrics. Following standard practices in zero-shot anomaly detection literature such as  [April_GAN, AnomalyCLIP], we evaluated our method using the AUROC. Additionally, we report APs for anomaly detection and use AUPRO [aupro] to assess pixel-level anomaly localization performance. Results are provided at both dataset-level and domain-level averages.
Table 1: Model performance on public benmarks. The top section reports image‑level results (AUROC, AP), and the bottom section reports pixel‑level results (AUROC, AUPRO). Best entries are marked with bold and second‑best entries with underline.
@@
Table 2: Comparison of PILOT vs. state‑of‑the‑art on industrial benchmarks. $\mathbf{Img}$ and $\mathbf{Pix}$ denote image‑level and pixel‑level metrics, respectively. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. Arrows indicate change in performance: $\uparrow$(increase >0.5>0.5), $\downarrow$(decrease <0.5<-0.5), and $\circ$(no significant change, $\pm 0.5$).
@@
3.1 Comparison with State‑of‑the‑Art
As reported in Table 1, PILOT establishes state-of-the-art performance across all benchmarks for both image-level anomaly detection and pixel-level localization in the zero-shot setting. This can be benefited from PILOT comprehensive approach in addressing critical challenges prevalent in ZSAD, that explicitly tackles domain shift in test phase. Prior approaches [saa, winclip] rely solely on predefined prompts, which may not generalize well to unseen domains, whereas methods such as [April_GAN, AdaCLIP, CoCoOp, AnomalyCLIP] employ lightweight visual adapters or a single learnable prompt to capture anomaly cues, but often overfit to the auxiliary training set, limiting robustness under distribution shift. PILOT distinguishes itself by dynamically identifying the most relevant prompts for each input and refining only those learnable prompt parameters during inference using high-confidence pseudo-labels from the unlabeled target data. Note that while existing baselines can utilize the proposed image-level TTA strategies, these often degrade pixel-level localization performance and do not yield consistent improvements across both detection and localization tasks. On the contrary, PILOT uniquely achieves gains in both image-level and pixel-level metrics under TTA, a phenomenon analyzed in detail in Section 3.2. We also compare against state‑of‑the‑art full‑shot methods in Appendix F.1, which show that PILOT achieves image‑level accuracy on par.
@@ Appendix A Related Work
Zero-shot Anomaly Detection. ZSAD has gained increasing attention as a solution for identifying anomalies in unseen categories without requiring task-specific training data [aota2023zero, winclip]. Leveraging the generalization capabilities of Vision–Language Models (VLMs) such as CLIP [clip], many approaches extend these models by linking image features to textual descriptions of normal and anomalous states. For instance, CLIP-AD [clipad] and ZOC [zoc] are early studies that leverage CLIP for ZSAD, but primarily focus on anomaly classification rather than localization. WinCLIP [winclip] uses multiple handcrafted prompts with iterative patch-based forward passes, and April-GAN [April_GAN] refines local semantics via fine-tuning on an auxiliary dataset; both methods address image-level anomaly detection and pixel-level localization. More recent prompt learning methods, such as AnomalyCLIP [AnomalyCLIP], AnomalyGPT [anomalygpt], and AdaCLIP [AdaCLIP], optimize object-agnostic text prompts in a single pass using auxiliary dataset. However, reliance on auxiliary datasets may introduce bias and compromise the intended zero-shot setting. This trade-off has been examined by Marzullo et al. [marzullo2024exploring], who observed that while avoiding auxiliary data reduces fine-tuning overhead, static prompts often fail to adapt to domain shifts; conversely, fine-tuning on auxiliary sets can improve accuracy but risks dataset bias and reduced generalization. Our work aims at addressing these challenges by introducing a dual-branch prompt learning mechanism that combines learnable and fixed prompts, enabling more robust and adaptive anomaly detection under domain shift.
@@
CLIP Adaptation. Recent advances in CLIP fine‑tuning primarily rely on prompt learning, in which learnable tokens are introduced into the text encoder to align visual features with related semantics [CoOp]. Approaches such as CoCoOp [CoCoOp] and visual prompt tuning [visualpromptting] condition the prompts on image content to improve robustness under domain shift. MaPLe [khattak2023maple] extends prompt learning by inserting learnable tokens into both the visual and text encoders of CLIP, jointly optimizing them to enhance cross-modal alignment. These techniques achieve parameter‑efficient adaptation by updating only the prompt representations and, in some cases, only CLIP’s final visual projection layer [fahes2024fine]. Following this rationale, our method fine‑tunes the visual projection layer to minimize additional parameters.
@@
Test Time Adaptation. A related direction for adapting CLIP-based models, particularly under domain shift, is TTA. In TTA, the model updates its parameters using unlabeled test data, often guided by pseudo-labels or self-supervised objectives. For instance, self-supervised losses based on the model’s own outputs can fine-tune representations without ground-truth labels [sun2020test], and momentum-based pseudo-labeling helps enforce consistency across different augmented views [goyal2022test]. Other TTA strategies include entropy minimization, which guides the model to assign higher confidence to its predictions on unlabeled data, and the online updating of batch normalization statistics, which helps the model better align with the target distribution during test time. These approaches have been shown to improve robustness and generalization in image classification [nado2020evaluating] and segmentation [colomer2023adapt]. However, recent studies have found that these techniques are often less effective when the domain shift is large or when the model relies on strong vision-language priors, as in the case of CLIP. To our knowledge, AnovL [deng2023anovl] is the only TTA method for zero-shot anomaly localization, which employs an adapter module for pixel-level refinement and thus introduces additional parameters while focusing solely on localization. In contrast, our method adaptively updates the relevant learnable prompts, incurring no additional parameters and delivering consistent improvements in both image-level anomaly detection and pixel-level anomaly localization.
@@ B.1 Benchmark Datasets
Appendix B provides a comprehensive review of the thirteen benchmark datasets utilized in our study, as summarized in Table 4. To rigorously evaluate both image-level detection and pixel-level localization capabilities of PILOT, we select six industrial inspection datasets. DAGM [dataset_DAGM] and DTD-Synthetic [dataset_DTD] comprise textured surfaces exhibiting subtle variations in material appearance, presenting unique challenges for texture-based anomaly detection due to the fine-grained and ambiguous nature of defects. In contrast, MVTec AD [dataset_mvtec], VisA [dataset_visa], BTAD [dataset_btad], and SDD [dataset_ksdd] focus on object-centric industrial scenarios, with defects such as cracks, scratches, dents, and missing components distributed across everyday manufactured and synthetic objects. These datasets are characterized by diverse object categories and high variability in anomaly manifestation, supporting comprehensive benchmarking of localization and generalization performance.
@@
To assess clinical applicability, we include seven medical imaging benchmarks. Three brain scan datasets, HeadCT [dataset_headCT], BrainMRI [dataset_brain], and Br35H [dataset_br35h], comprise both healthy and pathological subjects, facilitating binary anomaly classification tasks on volumetric neuroimaging data. These datasets encompass a spectrum of anatomical and pathological variations, thereby enabling evaluation of robustness to inter-patient heterogeneity and rare conditions. The remaining four datasets provide pixel-wise lesion annotations and consist exclusively of abnormal samples. ISIC [dataset_isic] features dermoscopic images annotated for skin lesions, while CVC-ClinicDB [dataset_ClinicDB], CVC-ColonDB [dataset_ColonDB], and Kvasir [dataset_Kvasir] contain endoscopic images of the gastrointestinal tract, highlighting a range of polyp and ulcer appearances. Since these medical localization datasets lack normal examples, test-time adaptation is not performed for them; their localization results without TTA are reported in Appendix G.
@@ B.2 Domain Discrepancy Analysis
To validate that MVTec AD and VisA indeed represent statistically distinct domains suitable for strict zero-shot evaluation, we quantitatively assessed the distributional discrepancy between their CLIP global feature embeddings. Specifically, we employed Maximum Mean Discrepancy (MMD2[gretton2012kernel], Fréchet Inception Distance (FID) [heusel2017gans], and the Kolmogorov–Smirnov (KS) [massey1951kolmogorov] test. Our analysis yielded $\mathrm{MMD}^{2}=0.0328$ ($p$-value =0.0099=0.0099), indicating a significant difference in embedding distributions. The FID score between the two datasets is $51.9$, further corroborating a substantial distributional gap. Additionally, the KS test computed across all feature dimensions revealed that the minimum, mean, and maximum $p$-values are 0, $5.16\times 10^{-5}$, and $0.0329$, respectively, suggesting statistically significant differences in nearly every dimension of the embedding space. These results collectively confirm that MVTec AD and VisA are non-overlapping in the learned representation space, thus justifying their use as auxiliary and target datasets in our zero-shot anomaly detection experiments.